OpenStudy (anonymous):

limit as x approaches -6 when f(x)=(21-x)^5/3

OpenStudy (fifciol):

just substitute -6 in place of x

OpenStudy (fifciol):

it's (21-(-6))^5/3=27^5/3

OpenStudy (fifciol):

then apply the property: \[c^{\frac{ a }{ b }}=\sqrt[b]{c^a}\]

OpenStudy (fifciol):

so it's \[\sqrt[3]{27^5}\]

OpenStudy (anonymous):

Oh, that makes alot of sense. Thanks

OpenStudy (fifciol):

so \[\sqrt[3]{27^3*27^2}=\sqrt[3]{27^3}*\sqrt[3]{729}=27*9=243\]